OpenStudy (ghostedly):

I think I have the right answer, just want to check if it's actually the answer... (-a)^12 / (-a)^7 Is the answer -a^5?

OpenStudy (anonymous):

no

jimthompson5910 (jim_thompson5910):

(-a)^12/(-a)^7 = (-a)^(12-7) = (-a)^5 (-a)^5 = (-a)*(-a)*(-a)*(-a)*(-a) (-a)^5 = (-1)*(-1)*(-1)*(-1)*(-1)*(a*a*a*a*a) (-a)^5 = -a^5 you have the right answer

OpenStudy (ghostedly):

Ahhh, thank you!

OpenStudy (anonymous):

sorry i got it wrong